Information

The two groups of workers have the follow labor demand and unionization relations:

  • Software Developers: LDSD = 600 – 15W, LSSD = -100 + 10W where W is equal to an hourly wage and L is in 10’s of workers.   The current Software Developer workforce includes a total of 1,800 highly skilled and productive workers.
  • Retail Clerks: LDRC = 120 – 12W, LSRC = -50 + 8W where W is equal to an hourly wage and L is in 1,000’s of workers. The current Retail Clerk workforce includes 18,000 moderately skilled and productive workers.

The goods market for labor unions in Utah is parameterized by QSU = -20 + 7p and QDU = 65 - P where Q is percentage of workers at subject firms becoming union members and P is union dues per week.

Question

Given what you understand about labor markets generally, the effect the formation of a union sector might have on a non-union sector, the sensitivity the state legislature, Governor’s Office of Economic Development, and Department of Workforce Services has with respect to unemployment, what would you recommend to this labor union and its organizers? Feel free to think this through and elaborate on this, but don’t let your response exceed 1,000 words. If you'd like to include economic models in your response it will be worth some meaningful extra credit for this module.
Answer #1

1) Given, LDSD = 600 – 15W, LSSD= - 100 + 10W

harmony wage and work will follow in this market when:

LDSD = LSSD

600-15W = - 100+10W

25W = 700 => W = 28

LDSD = 600-15(28) = 180 ; LSSD = - 100+10(28) = 180

in the event that Union is conceivable to get programming designers a raise of 10%, then this would raise compensation to [ 28 +28 of 10%] 30.8. In which case

LDSD = 600-15(30.8) = 138

that is work request would fall by 180-138 = 42 units of work who will get dislodged.

2) as discovered in the above question the association should guarantee a time-based compensation of 30.8 from 28.

3) LDRC = 120 – 12W, LSRC = - 50 + 8W

at eqm: LDRC = LSRC

120-12W = - 50+8W

20W = 170 => W = 8.5

LDRC = 120-12(8.5) = 18 ; LSRC = - 50 +8(8.5) = 18

on the off chance that the association conveys as guaranteed, a 10% raise in the present compensation would be 8.5+ 8.5 of 10% = 9.35

LDRC = 120-12(9.35) = 7.8 that is decrease sought after of 18-7.8 = 10.2 which is the quantity of retail representatives who will be dislodged.

4) As discovered in the section 3, the compensation would need to ascend to 9.35.

Much obliged to you